How to arrive at the answer to the following problem?
2 divided by the 6th root of 81.

Answer by Alan3354(69443)

How to arrive at the answer to the following problem?
2 divided by the 6th root of 81.
------------
It's good to eliminate radicals in the denominator, makes it much easier to add and subtract things. This is usually called "simplifying" the expression.
2/6th rt(81)
= 2/(cube root(9))
= 2*(cube root(81))/9
= 2*(cr(27) * cr(3))/9
= 2*3*(cr(3))/9
= 2cuberoot(3)/3
Using a calculator,
=~ 0.961499713